​A plant operation has fixed costs of $2,000,000 per year

A plant operation has fixed costs of $2,000,000 per year, and its output capacity is 100,000 electrical appliances per year. The variable cost is $40 per unit, and the product sells for $90 per unit. 


a. What is the break-even point in terms of units? 

b. What is the break-even point in terms of dollars? 

c. How much capacity has been used? 

d. If the variable cost is increased to $55, what is the percent reduction (or increase) to the profit? What is the break-even point in units now?
Answer #1

a. Break even point in terms of units = Fixed costs/ ( Selling price - Variable cost) = 2,000,000 / ( 90 -40) = 40,000 units

b. Break even point in terms of dollars = 40,000 * 90 = 3,600,000

c. Since nothing is explicitly mentioned, it is assumed that 100 % of the capacity has been used.

d. New break even point = Fixed costs/ ( Selling price - Variable cost) = 2,000,000 / ( 90 -55) = 57143 units

% reduction in profit = (3,000,000 - 1,500,000) / 3,000,000 = 50%

Profit at variable cost of $40 Profit at variable cost of $55
Sales 9,000,000 9,000,000
Less: Variable cost 4,000,000 5,500,000
Contribution margin 5,000,000 3,500,000
Less: Fixed costs 2,000,000 2,000,000
Profit 3,000,000 1,500,000